# DCA-Bench: A Benchmark for Dataset Curation Agents

### v1

@@ 3.4 Evaluation Pipeline
Note that this evaluation process has several intrinsic challenges. Firstly, due to the nuanced nature of text generation, different Curator outputs can refer to the same issue, while similar Curator outputs may also point to different issues. Fixed keyword [49] or rule-based code tests [12] are often insufficient to distinguish the nuances. Additionally, Curators may be able to identify the issues using only a portion of the contextual evidence annotated in the test case, which means that evaluation protocols requiring the inclusion of all the annotated contextual evidence can be overly stringent. While human experts can easily capture such nuances and properly evaluate the Curator performance, human evaluation is too expensive for a benchmark in practice. Recognizing that evaluating the Curator’s performance is easier than building the Curator agent, we propose an automatic evaluation pipeline where LLMs are leveraged to serve as the Evaluator.
@@
For each test case, the Evaluator receives the Curator’s output along with the hints and the ground truth annotations about the issues in the test case. Then, the Evaluator is asked to rate the Curator’s performance for each criterion with a real number from 0 to 1. The final score is a weighted sum of the three ratings, with weights of $0.85$, $0.15$, and $0.05$, reflecting their respective importance. The Evaluator then categorizes the Curator’s performance according to defined thresholds on the final score. More details about the weights and thresholds are provided in Appendix A.7.
To enhance the robustness of the evaluation, a simplified voting strategy similar to Verga et al. [50] was implemented to mitigate the influence of randomness. The strategy comprises three rounds of voting. In the first round, we collect results from $n$ Evaluators. If a consensus is reached, the decision is made with the consensus; however, if there is no consensus, we conduct another round of voting within $m$ Evaluators, and apply a majority vote to make the decision. In the event of a tie, additional votes will be conducted one by one until a definitive majority vote is determined. In practice, $m$ and $n$ are both set to 2.
@@
4 Experiments
In this section, we first verify the performance of the Evaluator, which we will show that it reliably aligns with human annotators. We then apply this Evaluator to test some baseline Curators on DCA-Bench and discuss the results.
4.1 Validation of the Evaluator
Experimental Setup
To verify the effectivene of the proposed Evaluator, we compare it with human annotations. We randomly select 23 test cases for each of the four hint levels, which results in 92 (issue, hint) pairs. We then collect the outputs of a baseline Curator agent, which is based on the OpenAI Assistant API with GPT-4-0125 preview equipped with the Code Interpreter tool [^4] See https://platform.openai.com/docs/assistants/tools/code-interpreter., on each of the 92 pairs. Then, we have the Evaluator and human annotators to rate the Curator outputs. For the Evaluator, we test four different models as the backend: GPT-4-0125 preview, GPT-4o, GPT-3.5-turbo, and Llama3-70B-Instruct. We use the chatCompletion API for OpenAI models and the Hugging Face interface for locally deploying Llama-3. For human annotators, we provide the codebook (see Appendix A.2) to follow. Finally, we compare the Evaluator’s ratings with the human annotations (ground truth), focusing on the binary classification between “fail” and (success/success+). For clarity, we refer to the latter as “succ” in the following context. The evaluation results using a three-class classification of “fail”, success, and success+ are provided in Appendix A.5.
Table 4: The performance of Evaluators with different backend models, using human annotations as ground truth. Here, we treat succ as the positive class when calculating Precision, Recall, and F1 Score. The “$\kappa$ Value” refers to Cohen’s $\kappa$ between the Evaluator ratings and the human annotations.
@@ 4.1 Validation of the Evaluator
GPT-4-0125-preivew 97.83 93.10 100.00 96.43 94.87
GPT-4o 92.39 81.25 96.30 88.14 82.59
GPT3.5-turbo 68.48 48.21 100.00 65.06 42.15
Llama3-70B-Instruct 65.22 45.61 96.30 61.90 36.69
Results
@@
As shown in Tab. 4, while the Evaluators with all backend models have high recall scores, which means they rarely misclassify succ as fail, it is quite often for weaker models (GPT3.5-turbo and Llama3-70B-Instruct) to misclassify fail as succ, as evidenced by the relatively low precision score.
Overall, the Evaluator with the backend GPT-4-0125-preview highly aligns with the human annotations, indicating that it could serve as a good proxy of human evaluations in our benchmark.
4.2 Benchmarking the Baseline Curators
Experimental Setup
We apply our Evaluator with backend GPT-4-0125-preview to benchmark the performance of some baseline Curators. We first experiment with the baseline Curator based on the OpenAI Assistant API equipped with the Code Interpreter tool on the full set of test cases and hint levels. As pointed out by OpenAI API users555see Section A.8 for details, the OpenAI Assistant API’s performance is often worse than that of the web-interface ChatGPT, so we also test Curators based on the web-interface ChatGPT. However, experimenting with the web-interface ChatGPT is not scalable, as we need to manually feed the prompts to the browser. Therefore, we selected 32 (issue, hint) pairs where the Assistant-API-based Curator fails, forming a hard set of DCA-Bench. Then, we carry out this small-scale experiment on ChatGPT-4, ChatGPT-4o, and ChatGPT-4 with reference materials. The model’s knowledge is limited to May 2024, when we conduct the experiments. The performance of the Curator is classified as fail and succ by the Evaluator.
Results
Fig. 4 shows the results of the Assistant-API-based Curator [^6] See Appendix A.9 for a cost analysis of evaluating baseline Curator.. Without any hints (Hint Level 0), the Curator only successfully detects 10.99% of issues. With more informative hints, the performance of the Curator increases accordingly, achieving a 62.64% success rate when given the most informative hint (Hint Level 3). Despite this improvement, the success rate is still unsatisfactory given the amount of information provided, as we cannot expect to have such informative hints when developing the real autonomous dataset curation pipeline.
Figure 4: Results of the baseline Curator based on the OpenAI Assistant API.
Table 5: Success rates of different models tested on the hard set of DCA-Bench. Even strong language models like ChatGPT-4 and ChatGPT-4o only succeeded in detecting approximately 19% of the issues. Interestingly, providing ChatGPT-4 with reference materials led to a drop in its performance.
Model Succ /%
ChatGPT-4 18.80
ChatGPT-4o 18.80
ChatGPT-4 w/refs 12.50
Tab. 4 shows the results of web-interface ChatGPT-based Curators on the hard set. As we can see, OpenAI’s most powerful models, ChatGPT-4 and ChatGPT-4o, only succeeded in detecting approximately 19% of the issues. Intuitively, dataset curation knowledge should help the assistant better identify the issue. Therefore, we used OpenAI’s MyGPT [^7] https://openai.com/index/introducing-gpts/ to add reference materials in DCA-Bench as knowledge to ChatGPT and tested its performance (the row ChatGPT-4 w/ refs). Surprisingly, the performance of the model drops in comparison to the model without such extra knowledge. We suspect that this is because the reference materials introduce too generic information, taking up the context window of the model and reducing its attention to the context relevant to the specific issues.
@@
Overall, the results of the baseline Curators suggest that the proposed DCA-Bench is a challenging benchmark for state-of-the-art LLMs.
5 Conclusion

### v2

@@ 2.4. Evaluation Pipeline
Note that this evaluation process has several intrinsic challenges. Firstly, due to the nuanced nature of text generation, different Curator outputs can refer to the same issue, while similar Curator outputs may also point to different issues. Fixed keyword (He et al., 2024a) or rule-based code tests (Jimenez et al., 2024a) are often insufficient to distinguish the nuances. Additionally, Curators may be able to identify the issues using only a portion of the contextual evidence annotated in the test case, which means that evaluation protocols requiring the inclusion of all the annotated contextual evidence can be overly stringent. While human experts can easily capture such nuances and properly evaluate the Curator performance, human evaluation is too expensive for a benchmark in practice. Recognizing that evaluating the Curator’s performance is easier than building the Curator agent, we propose an automatic evaluation pipeline where LLMs are leveraged to serve as the Evaluator.
@@
For each test case, the Evaluator receives the Curator’s output along with the hints and the ground truth annotations about the issues in the test case. Then, the Evaluator is asked to rate the Curator’s performance for each criterion with a real number from 0 to 1. The final score is a weighted sum of the three ratings, with weights of $0.85$, $0.15$, and $0.05$, reflecting their respective importance. The Evaluator then categorizes the Curator’s performance according to defined thresholds on the final score [^2] More details about the settings of weights and thresholds are provided in Appendix B.1..
While it’s natural to ask LLMs to directly classify the Curators’ performance into fail, success, success+, this strategy is poorly aligned with the annotation of human experts as we tested. In addition to common prompt methods (e.g., COT (Wei et al., 2023), role-playing (Kong et al., 2024), delimiter (OpenAI, 2024c)), we believe more prompt engineering is needed to improve the performance of the Evaluator, and here is our design:
We chose to have the Evaluator rate on specific metrics because this strategy allows for iterative refinement of the prompt by adjusting the weight and threshold of each metric. This approach decreases uncertainty by introducing fixed numerical comparisons, which provides a clearer direction for tuning, as opposed to relying solely on textual standards, which can be more ambiguous.
In designing the Evaluator’s prompt, we started by selecting a small subset of human-labeled samples. We began with an initial weight distribution of ($0.7,0.2,0.1$) to reflect the relative importance of each metric in Section 2.4. The high initial weight for ”Precise Contextual Evidence” was chosen to align with our focus on whether the Curator identifies annotated issues by providing contextual evidence, as discussed in Section 2.4. We then fine-tuned these weights using the selected subset to better align the Evaluator’s prompt with human judgments.
The motivation behind this design is to make iterative refinement of prompts more effective. Developing a good prompt requires testing it on a subset of human-labeled issues with corresponding outcomes from Curators and then making adjustments based on the outcomes. Unlike the model training process, which benefits from clear gradients and systematic optimization, tuning text prompts is often an empirical process necessitating numerous trials without the guarantee of improvement in its performance. Employing numerical weights and thresholds could enable us to tune the prompt more fine-grained, only changing the weights or threshold based on the results, which share similarities with linear regression, resulting in a more efficient prompt tuning process and better performance. By ”training” on a this human-labeled subset, we finalized weights in the prompt as $0.85$, $0.15$, and $0.05$.
To enhance the accuracy of the evaluation, a simplified voting strategy (Verga et al., 2024) was implemented to mitigate the influence of randomness. The strategy comprises three rounds of voting. In the first round, we collect results from $n$ Evaluators. If a consensus is reached, the decision is made with the consensus; however, if there is no consensus, we conduct another round of voting within $m$ Evaluators, and apply a majority vote to make the decision. In the event of a tie, additional votes will be conducted one by one until a definitive majority vote is determined. In practice, $m$ and $n$ are both set to 2.
@@
3. Experiments
In this section, we first assess the performance of the Evaluator, demonstrating its strong alignment with human annotations, along with its robustness and negligible bias. We then utilize this Evaluator to evaluate baseline Curators on DCA-Bench and further analyze the results.
3.1. Validation of the Evaluator
3.1.1. High Alignment with Human
To verify the effectiveness of the proposed agent Evaluator, we conduct a comparative analysis against human annotations. We randomly select 23 test cases for each of the four hint levels, resulting in a total of 92 (issue, hint) pairs. Using these pairs, we collect the outputs of a baseline Curator agent, which is based on the OpenAI Assistant API with GPT-4-0125-preview equipped with the Code Interpreter tool [^3] See https://platform.openai.com/docs/assistants/tools/code-interpreter., on each of the 92 pairs. The Evaluator and human annotators then rate the Curator outputs independently. To develop a reliable and cost-efficient Evaluator, we test following different models as backends: GPT-4o-2024-11-20, GPT-4o-2025-05-13 (OpenAI et al., 2024), o3-mini-2025-01-31 (OpenAI, 2024b), GPT-4-0125-preview (OpenAI, 2024a), GPT-3.5-Turbo (OpenAI, 2023a), Llama3.3-70B-Instruct (AI, 2024), Llama3-70B-Instruct(Dubey et al., 2024), Deepseek-R1 (DeepSeek-AI et al., 2025a), and Deepseek-V3 (DeepSeek-AI et al., 2025b). Details for setting up those Evaluators are included in Appendix B.2. Human annotators follow a structured codebook [^4] Please see Appendix B.4 for details to ensure rating consistency. Finally, we compare the Evaluator’s ratings with human annotations (ground truth) using a binary classification scheme: fail versus succ (where succ combines both success and success+). Additional results based on a three-class classification scheme (fail, success, and success+) are provided in Appendix B.5.
Table 3. The performance of Evaluators with different backend models, using human annotations as ground truth. We treat succ as the positive class when calculating Precision, Recall, and F1 Score. The “$\kappa$ Value” refers to Cohen’s $\kappa$ between the Evaluator ratings and the human annotations. Color Legend:   (95–100%),   (85–95%),   (60–85%),   (0–60%). Based on the table, gpt-4o-2024-11-20 and gpt-4-0125-preview outperform other models. Among open-source models, DeepSeek R1 and DeepSeek-V3 demonstrate reasonably good performance, though there remains a noticeable gap compared to the top-performing models. Conversely, LLama-3-70B-Instruct and GPT-3.5-Turbo exhibit the weakest performance. Given that gpt-4o-2024-11-20 is more cost-efficient than gpt-4-0125-preview, we primarily adopt it for evaluating Curators in Section 3.2.
@@ 3.1.1. High Alignment with Human
gpt-4o-2024-11-20 97.83 100.00 92.59 96.15 94.64
gpt-4-0125-preview 96.74 92.86 96.30 94.55 92.22
gpt-4o-2024-05-13 92.39 81.25 96.30 88.14 82.59
DeepSeek-R1 92.39 81.25 96.30 88.14 82.59
DeepSeek-V3 93.48 88.89 88.89 88.89 84.27
gpt-3.5-turbo 68.48 48.21 100.00 65.06 42.15
Meta-Llama-3-70B-Instruct 69.57 49.09 100.00 65.85 43.68
Meta-Llama-3.3-70B-Instruct 88.04 72.22 96.30 82.54 73.73
o3-mini-2025-01-31 91.30 100.00 70.37 82.61 77.04
@@
As shown in Tab. 3, Evaluators using all backend models achieve high recall scores, indicating that they rarely misclassify succ as fail. However, weaker models such as GPT-3.5-Turbo and Llama-3-70B-Instruct frequently misclassify fail as succ, as reflected in their relatively low precision scores.
Overall, Evaluators with the GPT-4o-2024-11-20 and GPT-4-0125-preview backends exhibit strong alignment with human annotations, suggesting that they can serve as reliable proxies for human evaluations in our benchmark. Other models, such as DeepSeek-R1 and DeepSeek-V3, demonstrate good performance, though a noticeable gap remains compared to the top two models.
3.1.2. Consistent Evaluation
To assess the consistency of ratings generated by our prompting and voting strategy, we conducted three independent evaluations using two Evaluators powered by GPT-4-0125-preview and GPT-4o-2024-11-20. Each Evaluator rated the same outputs produced by the GPT-4-0125-preview Curator on a subset of DCA-Bench (92 data-hint combinations). The results show high stability, with a standard deviation of ±$\pm 0.89\%$ for GPT-4-0125-preview and ±$\pm 0.51\%$ for GPT-4o-2024-11-20. Detailed numerical results are provided in Tab. 7 in the Appendix.
3.1.3. Negligible Self-Preference and Length Bias
To ensure the fairness of our evaluation strategy, we further examine two common biases that often arise when using LLMs as evaluators: self-preference bias and length bias. Through a thorough analysis, we find that our evaluation strategy remains robust against both biases.
Self-Preference Bias: We first assess under such designs, whether the Evaluator favors Curators that use the same model as their backend. Using GPT-3.5-Turbo and GPT-4-0125-preview as both Curators and Evaluators, we compute the Average Odds Difference (AOD) (Majumder et al., 2022), a fairness metric measuring disparities in true and false positive rates. The results indicate an AOD of $6.09\,(\pm 1.72)\%$ for GPT-3.5-Turbo and $5.31\,(\pm 2.76)\%$ for GPT-4-0125-preview, both well within the fairness threshold of $10\%$. Furthermore, GPT-4-0125-preview Evaluator aligns closely with human annotations across multiple evaluation metrics, confirming the negligible impact of self-preference bias.
Length Bias: To test whether response length influences evaluation outcomes, we modify Curator responses to be more verbose or succinct while preserving content. Experiments using GPT-3.5-Turbo-0125 and GPT-4o-mini-0718 show minimal differences in accuracy across different lengths (at most $\approx 1\%$ variation), indicating that response length does not significantly affect evaluations.
These findings confirm that our LLM Evaluator maintains fairness with negligible self-preference and length bias. Please see more details in Appendix B.7.
3.2. Benchmarking the Baseline Curators
3.2.1. Experimental Setup
In the following sections, we use the Evaluator powered by GPT-4o-2024-1120 due to its lower cost and strong performance, as shown in Tab.3. We also implement several baseline methods with different backend models, detailed in Tab.4.
Table 4. The performance of several Curators with different backend models across various hint levels on the full DCA-Bench dataset is evaluated. The Evaluator is based on the GPT-4o-2024-11-20 model. The success rate here refers to succ (success and success+, defined in Sec. 2.4). Notably, without any hints ($h_{0}$), even leading models such as DeepSeek-R1 and GPT-4o-2024-11-20 perform unsatisfactorily, detecting only about 30% of issues. As more informative hints are provided, the success rates increase accordingly. However, even with the most informative hints, the success rate of the models fails to exceed 80%. The results also indicate that incorporating RAG with knowledge relevant to dataset curation does not guarantee a performance improvement across all test cases and models.
Success Rate / %
Model Name $h_{0}$ $h_{1}$ $h_{2}$ $h_{3}$ Avg.
w/o Knowledge RAG
DeepSeek-R1 29.86 52.04 52.04 73.30 51.81
DeepSeek-V3 15.84 39.82 39.82 76.02 42.87
GPT-4-0125-preview 10.86 27.15 34.84 67.42 35.07
GPT-4o-2024-11-20 20.36 41.63 47.51 76.02 46.38
w/ Knowledge RAG
DeepSeek-R1 29.41 45.25 56.11 78.28 52.26
DeepSeek-V3 12.22 38.91 42.99 75.57 42.42
For open-source models, we utilize SGLang (Zheng et al., 2024) for local deployment and LangChain (Chase, 2022) to manage file-based context inputs. One intuitive way to improve the Curator’s performance is by incorporating external knowledge about dataset curation. To achieve this, we implement a RAG pipeline using LangChain. The RAG vector store is built using materials from DCA-Bench, as described in Section 2.1. Among commercial models, we access GPT-4-0125-preview via the OpenAI Assistant API, which includes the Code Interpreter tool. [^5] As noted by OpenAI API users, the performance of the OpenAI Assistant API can fluctuate. See Appendix E.1 for further details. For GPT-4o-2024-1120, we use the LangChain interface in the same way as for open-source models. Additional implementation details are provided in the Appendix B.3.
3.2.2. Results and Analysis
Tab. 4 presents the results of various baseline Curators evaluated using the Evaluator powered by the GPT-4o-2024-11-20 model. Among them, the Curator utilizing DeepSeek-R1 with RAG enabled achieves the highest average score of $53.05\%$, while GPT-4-0125-preview accessed via the OpenAI Assistant API performs the worst. Notably, without any hints ($h_{0}$), even the strongest models, such as DeepSeek-R1 and GPT-4o-2024-11-20, struggle to detect issues, achieving a success rate of only around 30%. As more informative hints are introduced, the success rates improve accordingly. However, even with the most informative hints, no model surpasses an 80% success rate.
We also observed that, although using RAG improves the performance of DeepSeek-R1, the improvement is not significant. Moreover, when applying the same RAG approach to the DeepSeek-V3 model, we even observed a decline in performance. Therefore, while leveraging RAG with dataset curation materials that provide curation knowledge is an intuitive approach, we believe it will need more designs and exploration to offer practical assistance to LLM agents in solving real-world dataset challenges. This could be attributed to several factors. We suspect that the reference materials introduce overly generic information, which occupies the model’s context window and diverts its attention from the specific issues at hand. Additionally, the diversity of these materials may not be sufficient to fully cover the range of issues present in DCA-Bench.
Thus, we believe that integrating more extensive and specialized knowledge, leveraging advanced RAG techniques, and fine-tuning models have great potential for improvement. Research on multi-agent collaborations (Wu et al., 2023a) also shows promising prospects. Further exploration is needed, and we leave it for future works.
Besides, we analyze the success rate of the Baseline Curator on issues with different token lengths. We observe that the success rate doesn’t monotonically decrease as content length grows. A detailed analysis of its performance across different issue types and hint levels is provided in Appendix C.
@@ 3.2.2. Results and Analysis
Overall, the results of the baseline Curators suggest that the proposed DCA-Bench benchmark presents a substantial and meaningful challenge for the AI community, highlighting the need for further exploration and research.
4. Ethics Statement
Our study involves human annotators, who are the authors themselves. The datasets utilized in DCA-Bench are sourced from various platforms, with proper licensing documentation provided. Additionally, we have compiled a detailed table (DCA_Benchmark.csv) in the supplementary materials, listing license information for all data files. Regarding secondary usage of user-generated data, we reviewed the terms of service for GitHub, HuggingFace, and Kaggle. Based on their policies and existing academic precedents, we believe our usage aligns with acceptable practices for non-commercial research purposes. We do not collect any user information and commit to removing content upon valid request.
Furthermore, our dataset includes 10 instances involving ethical or legal risks, with 2 data points potentially exhibiting bias. These cases are included to assess LLMs’ ability to identify and address ethical concerns, contributing to a realistic evaluation of AI safety. The annotation process poses minimal risk to annotators. Please see more details regarding the ethical considerations in Appendix D.
5. Related Work
Dataset Quality Management. With the advancement of modern AI, the need for high-quality datasets is surging (Jain et al., 2020; Kaplan et al., 2020). Consequently, open dataset platforms like Hugging Face and Kaggle keep growing rapidly. However, recent studies (Klie et al., 2024, 2023; Weber-Genzel et al., 2023) indicate that many popular datasets suffer from errors, biases, or annotation artifacts (Gururangan et al., 2018), and lack proper documentation (Yang et al., 2024). In addition, common issues in real-world datasets include problems with data loading scripts and file discrepancies, where conflicting information exists between files.
There have been many existing efforts on dataset quality management. The FAIR principles (Wilkinson et al., 2016) provide a broad framework for creating and managing datasets. Gebru et al. (2021) puts forward a standardized process for documenting datasets in the context of machine learning. Gong et al. (2023) propose several strategies, including dataset profiling, data cleansing, and quality monitoring. More recent works (Gebru et al., 2021; Wang et al., 2023; Gan et al., 2024) further explore standardizing dataset management practices. Based on these works, researchers have developed advanced dataset management toolkits and systems (Gupta et al., 2021; Mao et al., 2023; Zhou et al., 2024). More recently, Zhou et al. (2023); Chen et al. (2023b) developed systems that are capable of greatly enhancing data quality. However, they focus on filtering a large-scale LLM-pertaining corpus rather than detecting dataset quality issues. There are some specialized works focused on detecting harmful content (Althobaiti, 2022; Kirk et al., 2023), annotation errors (Weber-Genzel et al., 2023; Wang and Mueller, 2022) and data attributions (Longpre et al., 2023). Nevertheless, those methods are not generalizable enough and lack the flexibility to solve challenges such as nuanced ethical biases (see example A.1) and incorrect annotations involving logical or factual errors (see example A.3). Moreover, studies on file discrepancies (example A.2) and relevant issues remain scarce. Consequently, we still primarily rely on dataset users or platform maintainers to identify these issues in practice. We envision LLM agents as a promising technique to address the challenges in this space and propose a benchmark dataset as an initial step toward this goal.
LLMs for Software Engineering. Our work is also relevant to the emerging research direction on using LLMs to solve software engineering tasks. Fan et al. (2023) suggest that there are two major categories in this area: (i) code generation and completion, and (ii) maintenance and evolution.
There has been extensive literature on LLMs for code generation and completion, in terms of both models and benchmarks. On the model side, various code generation LLMs have been proposed (Chen et al., 2022; Luo et al., 2023; Rozière et al., 2024), along with agent frameworks designed to tackle complex coding problems, such as data analysis (Hong et al., 2024; Guo et al., 2024). On the benchmark side, datasets have been created to complete functions and short programs from natural language descriptions (Jimenez et al., 2024a; Chen et al., 2021; Austin et al., 2021; Lai et al., 2022; Zhang et al., 2024; Ding et al., 2023) and retrieve relevant code (Liu et al., 2023b). On the framework side, Reflexion (Shinn et al., 2023), Parsel (Zelikman et al., 2023) and MetaGPT (Hong et al., 2023) are the best reasoning frameworks at the time of writing. We note that the main purpose of the proposed DCA-Bench is to measure LLM agents’ capability of discovering hidden dataset issues instead of generating code, although generating and running some test code may be helpful.
For the second category, maintenance and evolution, it involves tasks such as localizing and fixing bugs (Wu et al., 2023b; Feng and Chen, 2023), improving program performance (Garg et al., 2024; Chen et al., 2023a), and refactoring code without changing program behavior (Poldrack et al., 2023). Our work falls into this category while differing from previous studies in two key aspects. Firstly, while most literature focuses on tasks with well-defined outcomes, such as fixing a buggy function, our work addresses non-standard issues like ethical concerns in data points or cross-file inconsistencies. Secondly, few studies have explored using LLMs to maintain dataset repositories, which is increasingly important for the machine learning community.
LLMs as A Proxy of Human Evaluation. Recent studies have shown that LLMs, when carefully prompted, can serve as a good proxy of human evaluations for a number of scenarios, such as evaluating text generation quality (Liu et al., 2023a; Chiang and Lee, 2023), reasoning ability (He et al., 2024b; Hao et al., 2024), and generated image quality (Chen et al., 2023c; You et al., 2024). It’s difficult to automatically evaluate the agents’ performance in complex generative tasks like text generation. Traditional reference-based metrics, such as BLEU and ROUGE, have been shown to have a relatively low correlation with human judgments (Liu et al., 2023a). These problems necessitate human intervention in evaluation judgment, making the process costly and hard to scale up. Facing these challenges, researchers have explored using LLMs for evaluations. Several existing works have shown that well-designed LLM evaluations align well with human evaluations through techniques including auto COT (Chiang and Lee, 2023), self-explaining rating (Chiang and yi Lee, 2023), and voting with multiple LLM evaluators (Verga et al., 2024). More and more researchers have applied LLMs in the evaluation process, including evaluating text generation quality (Liu et al., 2023a; Chiang and Lee, 2023), reasoning ability (He et al., 2024b; Hao et al., 2024), and generated image quality (Chen et al., 2023c; You et al., 2024).
6. Conclusion